[Brahms-l] Shift report 20050322 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Mar 22 2005 - 00:20:59 EST
Shifter leader was sjs
Shifters were 


Beam studies until 19:00.  Collision at 19:45, but had some trouble  
turning on D3.  
4deg A608; 40deg A700
Run      Trig 6      Comments
13906       90K      SMaxR trip
13907       64K      T3 trip, was tripped already for 13906, but 
                     not detected by BADTRIP. 
13908       251K     Good.  Ran for 1 hr.

4deg 1/2A;  40deg A700
Run       Trig 2     
13909       34.5K
13910       26.6K

Only detector problem noted was loss of wire 23 in T5-3-V2.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 22 00:21:16 2005

This archive was generated by hypermail 2.1.8 : Tue Mar 22 2005 - 00:21:29 EST